# Many Worlds in Theory Space: A Quantum Origin for the Constants of Nature

### v1

@@ Abstract
Modern physics relies on a specific set of dimensionless parameters—interaction strengths, mass ratios, and vacuum energy—whose observed values lie in extraordinarily narrow ranges required for complexity and life. Standard explanations often invoke external multiverses or ad hoc ensembles. This work proposes a purely quantum-mechanical origin for these constants by enlarging the configuration space of the Wheeler–DeWitt equation to include the theory-defining parameters themselves, and extends the Everett Many Worlds interpretation to include many worlds of different physical laws.
We treat fundamental constants not as fixed inputs but as dynamical quantum variables stabilized at the Grand Unified Theory (GUT) or Planck scale. The universal wavefunction thus carries amplitude over a “Theory Space” of possible physical laws. We demonstrate that early-universe symmetry breaking and decoherence naturally fragment this wavefunction into distinct, non-interacting “U-sectors,” each effectively a classical universe with fixed constants. In this extended Many-Worlds framework, fine-tuning is reinterpreted through quantum selection: while the vast majority of the cosmic wavefunction may describe sterile vacua, observers necessarily inhabit rare habitable branches. Finally, we suggest that the total integrated amplitude over theory space serves as a rigorous Bayesian evidence metric, enabling the comparative evaluation of candidate microscopic theories of quantum gravity.
@@ 1 Introduction
The “fine-tuning” of the universe refers to the empirical fact that the constants of nature, such as the fine-structure constant $\alpha$ or the cosmological constant $\Lambda$, occupy a tiny island of parameter space compatible with stable matter and long-lived stars [CarrRees1979, BarrowTipler1986].
The prevailing explanation for this puzzle relies on a synthesis of two conceptually distinct frameworks: the String Landscape and Eternal Inflation. It is useful to distinguish their roles:
@@
  • Eternal Inflation serves as the “engine.” It provides the dynamical mechanism to populate the landscape. Through continuous exponential expansion and bubble nucleation, it physically realizes the possibilities, creating vast, causally disconnected regions with different parameters.
Together, these form the Landscape Multiverse. Considerable theoretical effort has been directed toward understanding how these parameters could emerge from quantized degrees of freedom such as discrete fluxes and moduli fields. However, the Landscape Multiverse relies heavily on semiclassical reasoning; treating universes as bubbles nucleating in a background spacetime–and assumes the landscape structure a priori.
In contrast, the Many-Worlds Interpretation (MWI) of quantum mechanics [Everett1957] describes a multiverse generated internally by the unitary evolution of a single wavefunction. Standard MWI is limited, however: it generates branching histories (e.g., measurement outcomes) but not branching laws. All worlds in the standard MWI share the exact same electron mass and interaction strengths.
This paper unifies these perspectives by extending the canonical quantization of gravity. We propose that the “constants” of nature are quantum degrees of freedom that acquire fixed values through decoherence transitions in the very early universe. By enlarging the configuration space of the Wheeler–DeWitt equation [DeWitt1967], we construct a meta-wavefunction that superposes not just geometries, but entire theories.
@@
The remainder of the paper proceeds as follows. Section 2 details the high-energy origin of the constants. Section 3 constructs the enlarged configuration space. Section 4 describes the quantum structure of constants. Section 5 presents the meta-Wheeler–DeWitt equation.
@@ 2 The Quantum Origin of Physical Constants
A foundational principle of quantum mechanics is that all physical observables must correspond to linear operators acting on a Hilbert space. Yet, in the standard formulation of quantum field theory and cosmology, the physical constants of the Standard Model–gauge couplings, masses, and mixing angles–are not represented as operators but appear merely as fixed c-numbers in the Lagrangian. This indicates a conceptual incompleteness: the theory specifies a quantum dynamics for geometry and matter fields while treating the theory-defining structure itself as rigid classical background data.
@@
2.1 Spacetime Structure and Algebra
We treat the Planck-scale coherent state as the initial boundary condition for the effective theory. We remain agnostic regarding the mechanism that populated this state–whether via tunneling from nothing (nucleation), a quantum bounce, or an eternal quantum static state. Our focus is exclusively on the subsequent dynamical selection of physical laws via decoherence as the system evolves away from this highly symmetric limit.
The most fundamental constants are those that define the algebraic structure of the quantum spacetime manifold itself. In our framework, these are treated as the rigid scaffolding of the theory space, often set to unity in natural units, but physically distinct in their dimensional roles.
  • Constants: Speed of light ($c$), Planck constant ($h$), Boltzmann constant ($k$), Stefan-Boltzmann constant ($\sigma$).
  • Mechanism: These are fixed by the selection of the fundamental operator algebra and the macroscopic dimensionality $D$ of the uncompactified spacetime.
  • Physical Role: The speed of light $c$ defines the causal structure (Lorentz invariance) of the manifold, while $h$ sets the scale of quantum commutation relations.
@@ 2.4 Moduli Fields: Couplings and Mass Blueprints
  • Gravity ($G$): The gravitational constant is determined by the Volume Modulus. In string theory, the strength of gravity in 4D is inversely proportional to the volume of the compactified dimensions ($G_{4D}\propto G_{fund}/V_{compact}$).
@@ 2.5 Symmetry Breaking and Vacua
Setting Era Mechanism Constants Determineda Key Reference
String Scale
Stabilization of Radial Moduli;
Topology Change.
$G$ (Gravitational Const.)
$D$ (Dimensions)
Candelas et al. (1985) [Candelas1985]
Planck Era
Euclidean Path Integral;
Geometric Flux Cancellation.
$\Lambda$ (Vacuum Energy) Baum (1983) [Baum1983];
Hawking (1984) [Hawking1984]
GUT Scale
(Symmetry Breaking)
Flux Quantization;
Moduli VEVs (Dilaton);
Group Theory Factors.
$\alpha,\alpha_{s},e,\mu_{0},\epsilon_{0}$
$m_{e},m_{p},m_{n}$
(via Yukawas & $\Lambda_{\text{QCD}}$ seed)
Kaplunovsky (1988) [Kaplunovsky1988];
Susskind (2003) [Susskind2003]
  • a
    While masses ($m_{e},m_{p}$) manifest at low energy, their values are strictly determined by the dimensionless couplings (Yukawas and $\alpha_{s}$) fixed at the pre-inflationary GUT scale.
Table 1: The hierarchy of constants fixed in the pre-inflationary universe. All fundamental parameters are determined by geometric and topological transitions occurring before the onset of cosmic inflation, ensuring their universality across the observable cosmos.
@@ 3.1 Superspace and Theory Space
To describe a quantum origin for the laws of physics, we must treat these constants as coordinates. We introduce theory space ($\mathcal{T}$), a manifold where each point $\mathcal{T}=(D,X)$ represents a complete, self-consistent set of physical laws derived from a fundamental theory with total dimensionality $D_{\text{fund}}$.
@@ 4.1 U-Sector Hilbert Spaces and the Direct-Sum Structure
We acknowledge that we have no way of knowing how nature physically “stores” or operates on the wavefunction of the universe. The direct-sum formalism presented here is a mathematical suggestion–a rigorous way to describe a state space where the laws of physics themselves are quantum variables, while respecting the fact that observers within the system perceive those laws as fixed classical backgrounds.
@@ 4.2 The Extended Path Integral
To formalize this, we extend the gravitational path integral to include the theory space variables $\mathcal{T}\equiv(D,X)$ on the same footing as the geometric and matter degrees of freedom. Wheeler’s original spacetime foam concerned fluctuations in geometry at every point in space. We extend this to a global fundamental foam, where the constants themselves fluctuate. The extended functional integral becomes:
$\mathcal{Z}_{\rm ext}\;=\;\int\mathcal{D}_{\mathcal{T}}\mathcal{D}_{g}\mathcal{D}_{\phi}\;e^{\left\{i\,S[g,\phi,\mathcal{T}]\right\}}.$ (5)
@@
Here, the fundamental parameters—such as the effective gravitational constant $G(X)$, the cosmological constant $\Lambda(X)$, and the gauge couplings $g_{i}(X)$ within the matter Lagrangian—are determined by the moduli $X$.
Here, the fundamental parameters–such as the effective gravitational constant $G(X)$, the cosmological constant $\Lambda(X)$, and the gauge couplings $g_{i}(X)$ within the matter Lagrangian–are not constants but functions of the theory-space coordinates $X$. Here, the integration measure is nested: the global integration $\mathcal{D}_{\mathcal{T}}$ sums over all possible U-sectors (the choice of laws), while the inner functional integrations $\mathcal{D}_{g}\mathcal{D}_{\phi}$ sum over all local histories within that specific sector. $\mathcal{Z}_{\rm ext}$ thus represents the total amplitude of the cosmic landscape.
@@ 5.2 Generalization to Theory Space
The third term, $\hat{\mathcal{H}}_{\rm theory}$, is new. It governs the dynamics of the constants themselves—the transitions between different values of $D$ and $X$. It typically takes the form:
@@ 7.4 Sector Selection via Freezing
$\Psi(a\to\infty)\approx\sum_{i}C_{i}\,\Psi_{\text{U-sector }i}(a,\phi).$ (21)
@@ 8.1 The Peak Prediction: A Testable Constraint
Although the configuration space that allows complexity ($\Omega_{\rm obs}\subset\mathcal{T}$) is small relative to the huge space of possibilities, it occupies a significant volume. The quantum weighting framework predicts that the observed values of physical constants is most likely to sit near the maximum of the probability distribution, $P(\mathcal{T})$, within this habitable zone. This is the $\max P(\mathcal{T})$ prediction. If the observed universe were found in a low-probability tail region of $\Omega_{\rm obs}$, it would suggest the existence of a secondary, unknown filter beyond the initial quantum weighting. We are currently limited from testing this prediction because the volume and structure of the complexity-supporting region $\Omega_{\rm obs}$ have not been fully mapped; however, future computational work may render this constraint testable.
8.2 A Criterion for Fundamental Theories
Beyond explaining fine-tuning, the extended path integral formulation offers a quantitative tool for comparing candidate microscopic theories. In Bayesian inference, the relative plausibility of two models, $\mathcal{M}_{A}$ and $\mathcal{M}_{B}$, given observational data $D_{\rm obs}$ (i.e., the measured constants), is determined by the ratio of their evidences (the Bayes Factor):
$\frac{P(\mathcal{M}_{A}|D_{\rm obs})}{P(\mathcal{M}_{B}|D_{\rm obs})}=\frac{P(D_{\rm obs}|\mathcal{M}_{A})}{P(D_{\rm obs}|\mathcal{M}_{B})}\cdot\frac{P(\mathcal{M}_{A})}{P(\mathcal{M}_{B})}.$ (23)
In our framework, the “evidence” $P(D_{\rm obs}|\mathcal{M})$ corresponds to the total quantum measure of the U-sectors that are compatible with observation. If $\Omega_{\rm obs}\subset\mathcal{T}$ is the region of theory space containing constants compatible with our existence, then the evidence provided by a specific fundamental theory is:
$E(\mathcal{M})\;=\;\int_{\Omega_{\rm obs}}\mathcal{D}_{T}^{\mathcal{M}}\left|Z(T)\right|^{2}.$ (24)
This functional allows us to rate competing hypotheses about high-energy physics (e.g., different compactification manifolds or flux stabilization mechanisms). A theory that naturally generates a “fertile” landscape–one where $\Omega_{\rm obs}$ occupies a large volume of the configuration space–will yield a larger $E(\mathcal{M})$ and is statistically favored over a theory that requires extreme fine-tuning to produce a single viable U-sector.
Thus, $\mathcal{Z}_{\rm ext}$ transforms the measure problem from an interpretational difficulty into a rigorous selection criterion for ultraviolet completions of the Standard Model.

### v2

@@ Abstract
Modern physics relies on a specific set of parameters, interaction strengths, mass ratios, and vacuum energy. The observed values of these lie in a range extremely small compared to the required for complexity and life. Standard explanations often invoke external multiverses or ad hoc ensembles. This work proposes a purely quantum-mechanical origin for these constants. We do this by enlarging the configuration space of the Wheeler–DeWitt equation to include the theory-defining parameters themselves. We also extend the Everett Many Worlds interpretation to encompass many worlds governed by different physical laws.
Building on the anthropic framework originally suggested by Carr and Rees (1979), this work develops a rigorous quantum-mechanical origin for these constants by enlarging the configuration space of the Wheeler–DeWitt equation. We treat fundamental constants not as fixed inputs but as dynamical quantum variables stabilized at the Grand Unified Theory (GUT) or Planck scale. We show that early-universe symmetry breaking and decoherence naturally fragment this wavefunction into distinct universe sectors (U-sectors) in Hilbert Space. Each is effectively a classical universe with fixed constants. In this extended Many-Worlds framework, fine-tuning is reinterpreted through quantum selection: while the vast majority of the cosmic wavefunction may describe sterile vacua, observers necessarily inhabit rare habitable branches.
@@ 1 Introduction
The “fine-tuning” of the universe refers to the empirical fact that the constants of nature—such as the fine-structure constant $\alpha$ or the cosmological constant $\Lambda$—occupy a tiny island of parameter space compatible with stable matter and long-lived stars. This puzzle was famously highlighted by Carr and Rees [CarrRees1979], who presciently suggested that the Many-Worlds Interpretation of quantum mechanics might provide a natural framework for an ensemble of universes with varying constants.
Currently, the prevailing explanation for this puzzle relies on a synthesis of the String Landscape and Eternal Inflation. It is useful to distinguish their roles:
@@
  • Eternal Inflation serves as the “engine.” It provides the dynamical mechanism to populate the landscape. Through continuous exponential expansion and bubble nucleation, it physically realizes the possibilities on the menu, creating vast, causally disconnected regions with different parameters.
Together, these form the Landscape Multiverse. Considerable theoretical effort has been directed toward understanding how these parameters could emerge from quantized degrees of freedom—such as discrete fluxes and moduli fields—within this framework. However, the Landscape Multiverse relies heavily on semiclassical reasoning—treating universes as bubbles nucleating in a background spacetime—and assumes the landscape structure a priori.
This paper unifies these perspectives by extending the canonical quantization of gravity to realize the quantum ensemble required to resolve the fine-tuning paradox. We propose that the “constants” of nature are quantum degrees of freedom that acquire fixed values through decoherence transitions in the very early universe. By enlarging the configuration space of the Wheeler–DeWitt equation [DeWitt1967], we construct a meta-wavefunction that superposes not just geometries, but entire theories.
@@
The remainder of the paper proceeds as follows. Section 2 classifies the high-energy mechanisms—such as moduli stabilization and flux quantization—that determine the physical constants. Section 3 constructs the enlarged configuration space, treating these parameters as coordinates in a fiber bundle. Section 4 defines the Hilbert space structure of the U-sectors. Section 5 presents the meta-Wheeler–DeWitt equation and the Hamiltonian dynamics of the theory parameters during the coherent epoch. Section 7 traces the cosmological evolution of the wavefunction, deriving the selection of a specific U-sector via the freezing of moduli fields. Section 6 discusses the emergence of superselection rules that isolate these sectors. Finally, Section 8 addresses causality constraints, the derivation of the Standard Model Hamiltonian from the eternal framework, and the utility of this model for Bayesian theory comparison, followed by conclusions in Section 9.
@@ 2 The Quantum Origin of Physical Constants
A foundational principle of quantum mechanics is that all physical observables must correspond to linear operators acting on a Hilbert space. Yet, in the standard formulation of quantum field theory and cosmology, the physical constants of the Standard Model, gauge couplings, masses, and mixing angles, are not represented as operators but appear merely as fixed c-numbers in the Lagrangian. This indicates a conceptual incompleteness: the theory specifies a quantum dynamics for geometry and matter fields while treating the theory-defining structure itself as rigid classical background.
@@
2.1 Spacetime Structure and Units
The first class of constants consists of those that define the fundamental scales of the spacetime manifold. In our framework, these are treated as the fixed background units of the theory space. It is standard practice to normalize these parameters to unity (Planck units), revealing that they do not represent dynamical degrees of freedom but rather the dimensional scaffolding against which other variables are measured. These constants set the conversion factors between space, time, energy, and temperature. Other constants, such as the Stefan-Boltzmann constant $\sigma$, are derived directly from these fundamental units and the dimensionality $D$:
$\sigma=\frac{\pi^{2}k_{B}^{4}}{60\hbar^{3}c^{2}}\quad(\text{for }D=4).$ (1)
Since their values can be set to 1 by a choice of units, they do not constitute ”fine-tuning” in the same sense as the dimensionless couplings; rather, they define the scale of the arena in which the tuning occurs.
@@ 2.4 Moduli Fields: Couplings and Mass Blueprints
  • Gravity ($G$): The gravitational constant is determined by the Volume Modulus. In string theory, the strength of gravity in 4D is inversely proportional to the volume of the compactified dimensions ($G_{4D}\propto G_{\rm fund}/V_{compact}$).
@@ 2.5 Symmetry Breaking and Vacua
Class Constants Fixed GUT/Planck Mechanism
Spacetime
Topology
$D_{\text{fund}},D,d,c,h,k,\sigma$ Quantized Dimensionality: The total dimension $D_{\text{fund}}$ is fixed by quantum consistency (e.g., anomalies). The macroscopic $D$ is fixed by topological compactification.
Unified
Couplings
$G,\alpha,\alpha_{W},\alpha_{s},e,\epsilon_{0},\mu_{0}$ Moduli Stabilization: Scalar fields (Dilaton/Volume moduli) settle into potential minima, fixing the strength of gravity and the unified gauge forces ($SU(3){\times}SU(2){\times}U(1)$).
Mass
Generation
$m_{e},m_{p},R_{\infty},M_{W}$ Blueprint Fixing: The high-energy moduli fix the Yukawa couplings and Higgs potential parameters. This pre-determines the masses ($m\sim y\cdot v$) and confinement scales ($\Lambda_{QCD}$) that emerge at lower energies.
Vacuum
Energy
$\Lambda$ (Cosmological Constant) Flux Summation: Discrete fluxes and non-perturbative effects generate the vacuum energy density, which is fine-tuned by the vast number of flux choices (the Landscape).
Table 1: Classification of Physical Constants by GUT-Scale Origin. All fundamental parameters are determined by geometric and topological transitions occurring before the onset of cosmic inflation, ensuring their universality across the observable cosmos.
@@ 3.1 Superspace and Theory Space
To describe a quantum origin for the laws of physics, we must treat these constants as coordinates. We introduce theory space ($\mathcal{T}$), a manifold where each point $\mathcal{T}=(D,X)$ represents a complete, self-consistent set of physical laws derived from a fundamental theory with total spacial dimensionality $D_{\text{fund}}$.
@@ 4.1 U-Sector Hilbert Spaces and the Direct-Sum Structure
We acknowledge that we have no way of knowing how nature physically “stores” or operates on the wavefunction of the universe. The direct-sum formalism presented here is a mathematical suggestion; a rigorous way to describe a state space where the laws of physics themselves are quantum variables, while respecting the fact that observers within the system perceive those laws as fixed classical backgrounds.
@@ 4.2 The Extended Path Integral
To formalize this, we extend the gravitational path integral to include the theory space variables $\mathcal{T}\equiv(D,X)$ on the same footing as the geometric and matter degrees of freedom. Wheeler’s original spacetime foam concerned fluctuations in geometry at every point in space. We extend this to a global fundamental foam, where the constants themselves fluctuate. The extended functional integral becomes: The extended spacetime–foam functional integral can then be written as
$\mathcal{Z}_{\rm ext}\;=\;\int_{\mathcal{T}}\mathrm{d}\mu_{\mathcal{T}}(T)\int\mathcal{D}_{g}\,\mathcal{D}_{\phi}\;e^{iS[g,\phi;T]}.$ (6)
Here $\mathrm{d}\mu_{\mathcal{T}}(T)$ denotes the measure on theory space $\mathcal{T}$, while $\mathcal{D}_{g}$ and $\mathcal{D}_{\phi}$ are the usual functional measures over metrics and matter fields for fixed $T$. The theory–space measure may be written schematically as
$\mathrm{d}\mu_{\mathcal{T}}(T)\;=\;\sum_{D}\left[\sqrt{\det\mathcal{G}_{AB}(X)}\,\mathrm{d}^{n_{D}}X\right],$ (7)
where the sum runs over the discrete macroscopic dimensionalities $D$, the coordinates $X^{A}$ parametrize the continuous moduli and flux data for fixed $D$, $\mathcal{G}_{AB}(X)$ is the metric on that moduli submanifold, and $n_{D}$ is its dimension. Thus $\mathrm{d}\mu_{\mathcal{T}}(T)$ compactly represents the combined sum over discrete labels and integral over continuous theory–space coordinates.
@@
Here, the fundamental parameters, such as the effective gravitational constant $G(X)$, the cosmological constant $\Lambda(X)$, and the gauge couplings $g_{i}(X)$ within the matter Lagrangian, are determined by the moduli $X$.
Here, the fundamental parameters, such as the effective gravitational constant $G(X)$, the cosmological constant $\Lambda(X)$, and the gauge couplings $g_{i}(X)$ within the matter Lagrangian, are not constants but functions of the theory-space coordinates $X$. Here, the integration measure is nested: the global integration $\mathcal{D}_{\mathcal{T}}$ sums over all possible U-sectors (the choice of laws), while the inner functional integrations $\mathcal{D}_{g}\mathcal{D}_{\phi}$ sum over all local histories within that specific sector. $\mathcal{Z}_{\rm ext}$ thus represents the total amplitude of the cosmic landscape.
@@ 5.2 Generalization to Theory Space
The third term, $\hat{\mathcal{H}}_{\rm theory}$, is new. It governs the dynamics of the constants themselves, the transitions between different values of $D$ and $X$. It typically takes the form:
@@ 7 Cosmological Evolution of the Meta-Wavefunction
7.1 The Inertia of Physical Laws
To understand why the constants of nature appear fixed today, we must examine the kinetic energy associated with their variation. We start with the kinetic part of the Action ($S_{\rm kin}$) for the theory-space variables. In a spacetime with metric $g_{\mu\nu}$, the kinetic action for a dimensionless modulus $\chi$ takes the form:
$S_{\rm kin}=\int d^{4}x\sqrt{-g}\frac{f(\chi)}{2}(\partial_{\mu}\chi)^{2}\approx\int dt\,\underbrace{\frac{1}{2}\left(a(t)^{3}M_{\rm fund}^{2}\right)\dot{\chi}^{2}}_{L_{\rm kin}}.$ (16)
The term $f(\chi)$ is the Field Space Metric (or Kinetic Function) for the modulus field $\chi$. Here, the coefficient of $\dot{\chi}^{2}$ acts as the effective mass or inertia of the variable $\chi$:
$M_{\rm eff}(a)\equiv a(t)^{3}M_{\rm fund}^{2}.$ (17)
This inertia determines the magnitude of the Hamiltonian kinetic energy operator, $\hat{K}_{\chi}=\frac{\hat{P}_{\chi}^{2}}{2M_{\rm eff}}$.
  • In the Planck Era ($a\sim l_{P}$), the inertia is negligible ($M_{\rm eff}\sim 1$). The Hamiltonian kinetic term $\hat{K}_{\chi}$ is large, allowing $\chi$ to fluctuate wildly.
  • In the Macroscopic Era ($a\to\infty$), the inertia becomes astronomical. The Hamiltonian kinetic term is suppressed by a factor of $1/a^{3}$, effectively freezing the laws of physics in place.
@@ 7.5 Sector Selection via Freezing
$\lim_{a\to\infty}\Psi(a,\phi)\approx\sum_{i}C_{i}\,\Psi_{\mathrm{U}_{i}}(a,\phi).$ (25)
@@ 8.1 The Peak Prediction: A Testable Constraint
The quantum weighting framework predicts that the observed values of physical constants is most likely to sit near the maximum of the probability distribution, $p(\mathcal{T})$, within this habitable zone. This is the $\max p(\mathcal{T})$ prediction. If the observed universe were found in a low-probability tail region of $\Omega_{\rm obs}$, it would suggest the base theory is somehow deficient. We are currently limited from testing this prediction because the volume and structure of the complexity-supporting region $\Omega_{\rm obs}$ have not been fully mapped; however, future computational work may render this constraint testable.
8.1.1 Anthropic conditioning
Beyond explaining fine-tuning, the extended path integral formulation offers a quantitative tool for comparing candidate combinations of local and global theories. In the extended path-integral framework, a field configuration $\mathfrak{T}$ induces a quantum measure on theory space $\mathcal{T}$. For fixed laws $T\in\mathcal{T}$, we define the sectoral path integral
$\mathcal{Z}_{\rm ext}\;=\;\int\mathrm{d}\mu_{\mathcal{T}}(T)\;\mathcal{D}g\,\mathcal{D}\phi\;e^{\,i\,S[g,\phi;T]}.$ (27)
and the corresponding unnormalized weight
$w_{\mathfrak{T}}(T)\;=\;\bigl|Z_{\mathfrak{T}}(T)\bigr|^{2}.$ (28)
In practice, all known mechanisms for chemistry, stable atoms, and long-lived stars require three large spatial dimensions. We therefore restrict the anthropically relevant regions $\Omega_{\rm comp}$ and $\Omega_{\rm obs}$ to the $D=3$ slice of theory space, $\mathcal{T}_{3}\subset\mathcal{T}$, and normalize all conditional probabilities over this subspace:
$p_{\mathfrak{T}}(T\mid D=3)\;=\;\frac{\bigl|Z_{\mathfrak{T}}(T)\bigr|^{2}}{\displaystyle\int_{\mathcal{T}_{3}}\bigl|Z_{\mathfrak{T}}(T^{\prime})\bigr|^{2}\,\mathrm{d}\mu_{\mathcal{T}}(T^{\prime})},\qquad T\in\mathcal{T}_{3}.$ (29)
Let $\Omega_{\rm comp}\subset\mathcal{T}$ denote the complexity–permitting region of theory space: the set of U-sectors whose constants lead to long–lived stars, heavy elements, and nontrivial chemistry. Our actual observational situation (some complex observers exist) implies that we are necessarily located in $\Omega_{\rm comp}$, never in its complement. Anthropic reasoning corresponds to conditioning $p_{\mathfrak{T}}(T)$ on the event “complexity exists”:
$p_{\mathfrak{T}}\bigl(T\,\big|\,\mathrm{complexity}\bigr)\;=\;\frac{p_{\mathfrak{T}}(T\mid D=3)\;\mathbf{1}_{\Omega_{\rm comp}}(T)}{\displaystyle\int_{\Omega_{\rm comp}}\mathrm{d}\mu_{\mathcal{T}}(T^{\prime})\;p_{\mathfrak{T}}(T^{\prime})},$ (30)
where $\mathbf{1}_{\Omega_{\rm comp}}$ is the indicator function of the complexity–permitting region.
Our empirical data are more specific than the mere existence of complexity: we observe a definite set of constants. Let $\Omega_{\rm obs}\subset\Omega_{\rm comp}$ be the subset of theory space whose constants are compatible (within experimental tolerances) with the values we actually measure in our universe. The evidence for a given microscopic theory $\mathfrak{T}$, given these observations, is then
$E(\mathfrak{T})\;=\;\int_{\Omega_{\rm obs}}\mathrm{d}\mu_{\mathcal{T}}(T)\;p_{\mathfrak{T}}\bigl(T\,\big|\,\mathrm{complexity}\bigr).$ (31)
If $\Omega_{\rm obs}$ is small compared with the scales on which $p_{\mathfrak{T}}$ varies, this reduces to
$E(\mathfrak{T})\;\propto\;p_{\mathfrak{T}}\bigl(T_{\rm obs}\,\big|\,\mathrm{complexity}\bigr),$ (32)
where $T_{\rm obs}$ is a representative point in $\Omega_{\rm obs}$. [^2] For model comparison, only ratios of evidences matter; an overall proportionality constant cancels.
In this formulation, different UV completions of gravity and particle physics are not merely required to produce some complex U-sector. Rather, they are distinguished by how much conditional quantum weight they assign to the small region $\Omega_{\rm obs}$ within the full complexity–permitting domain $\Omega_{\rm comp}$. A theory for which our observed constants are typical among its complexity-supporting sectors (large $E(\mathfrak{T})$) is favored over a theory that produces our constants only in an exponentially suppressed corner of its theory space.